1,400 tons of wheat of two different varieties was delivered to the silo. When processing one variety, there was 2% of waste and 3% during the processing of the second variety of wheat. After the processing, the amount of remaining wheat was 1,364 tons. How many tons of each variety of wheat was delivered to the silo?
Answer by ikleyn(52878)

.
1,400 tons of wheat of two different varieties was delivered to the silo.
When processing one variety, there was 2% of waste
and 3% during the processing of the second variety of wheat.
After the processing, the amount of remaining wheat was 1,364 tons.
How many tons of each variety of wheat was delivered to the silo?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    x +     y = 1400          (1)
0.02x + 0.03y = 1400-1364.    (2)

Simplify:

    x +     y = 1400          (1')
0.02x + 0.03y =   36.         (2')

Multiply the first eq. by 2. Multiply the second eq. by 100. You will get

2x +  y = 2800,               (1'')
2x + 3y = 3600.               (2'')

Subtract eq. ('1'') from the eq. (2''). You will get

2y = 800  ---->  y = 400.

Then from (1) x = 1000.

Answer. 1000 tons of the first variety and 400 tons of the second variety.

This problem came to the forum earlier. I solved it in
https://www.algebra.com/algebra/homework/percentage/percentage.faq.question.1065528.html